For a binomial distribution, prove that the variance is less than the mean.
1
Expert's answer
2020-12-09T03:35:26-0500

Let X be a binomial variate with parameters n and p. Then,

Mean = np

And, Variance = npq

Mean - Variance = np−npq=np(1−q)=np2

Mean−Variance>0     [n∈N,p>0,therefore,np2>0]

Mean>Variance